# Weaver: Interweaving SQL and LLM for Table Reasoning

### v1

@@ Appendix B Dataset Details
- Short-Form Answering (WikiTQ): WikiTQ is a dataset designed for short-form answering where the steps to reach the answer can be relatively complex and the expected answer is a short piece of information. For example, the query "Which country had the most competitors?" in Figure 1. This task is ideal for testing how well information retrieval from a semi-structured table can be handled.
- Fact-Checking (TabFact): TabFact focuses on fact-checking tasks where the query involves verifying whether a particular statement is true or false. For example, the query, "Is the GDP of Japan in 2022 greater than that of Germany?" evaluates the framework’s ability to correctly interpret data points in complex tables and make valid judgments.
- Numerical Reasoning (FinQA): FinQA (Financial QA) is a dataset that requires the model to perform arithmetic operations or infer relationships between numerical values across different columns. For example, "What is the total revenue of Company X in 2021 after deducting expenses?" tests the model’s ability to handle numerical data and apply operations such as summation, subtraction, or other mathematical reasoning tasks.
@@
${}_{\text{MM}}$ is a multimodal dataset that requires reasoning across both tabular data and visual information. Unlike traditional table QA tasks, it incorporates both text and images within its tables, challenging the model to integrate structured (tables), unstructured (text), and visual (images) modalities. The data set consists of query demands that combine textual descriptions, numerical data, and visual cues from images. For example, a question might ask about the relationship between financial data in the table and trends depicted in an image. ${}_{\text{MM}}$ tests the model’s ability to perform complex multimodal reasoning, requiring SQL, LLM, and VLM calls to accurately derive answers. This makes it a valuable benchmark for evaluating systems that integrate and reason over multimodal inputs.
@@ Appendix C Few-Shot Prompting for Plan Generation
We use a few-shot prompting approach to generate plans during the planning stage. Through our analysis of various complex data problems, we identified three broad categories of transformation challenges that commonly arise. For each category, we manually crafted a representative example—rather than using examples from any particular dataset—to serve as in-context prompts. This was done to avoid memorization bias and to encourage cross-dataset generalization. These examples were specifically designed to capture the reasoning skills required for hybrid queries. We will include them in the appendix for clarity and reproducibility in the final version. Below are the three categories and the corresponding examples with sample tables:
@@
Listing 19: Few-shot examples for semantic reasoning from textual content
@@
"Indulge your senses with this botanical blend of rosemary and lavender. Gently cleanses while nourishing your hair, leaving it soft, shiny, and revitalized." 7.99 4.99
@@
Listing 20: Few-shot examples for commonsense or background knowledge inference
@@
Question: Which location had the maximum time taken between dispatch at one location and arrival or delivery at a subsequent location?
@@
Step 3: SQL - Pair Dispatched events with the corresponding Arrived at Hub or Delivered events for each order/product and calculate the time difference.
Step 4: SQL - Display the final output with paired events, durations, and relevant information sorted by Order_ID and Product.
@@
Listing 21: Few-shot examples for pre-processing or normalization of non-SQL-friendly columns
@@
Step 1: LLM - Format the column Placing by extracting only numerical values (e.g. 5 from Semifinal (5th)) and converting the text into numbers (e.g. Semifinal to 5, 12th to 12, 19 to 19).
Step 2: SQL - Retrieve the highest placing (rank) from the placing column by selecting the minimum number in the list as lower number corresponds to higher rank.
Appendix D Additonal Discussion
@@
Our core contribution lies in how we effectively leverage LLMs beyond what traditional SQL systems can offer, specifically in tasks involving column transformation and semantic inference. Weaver integrates LLM reasoning for two primary purposes: (i) handling tasks that SQL cannot perform—such as entity extraction or parsing ambiguous formats A.2 and (ii) inferring implicit knowledge based on pretraining, where the LLM is tasked with verifying a plan that involves semantic understanding Figure 2.
A detailed example from the WikiTQ dataset in Appendix A.2(Listing 12) demonstrates the LLM’s capability to normalize complex date strings like "Spring 1932" or "1935/36" into standardized YYYY formats. This transformation, impossible through SQL alone, was completed through a single LLM step with high consistency.
Listing 22: Example explaination
@@
Listing 23: Example explaination
Question: "What movies suitable for kids with positive reviews should be recommended based on their reviews after 2018?"
@@
LLM: Utilize a Large Language Model (LLM) to evaluate whether a movie is suitable for children by processing:
@@
Beyond Semi-structured Text.
We define semi-structured tables following prior work Gupta et al. (2020), focusing on unstructured or free-form content embedded within structured table formats—such as textual cells requiring semantic interpretation—rather than on hierarchical structures like JSON or HTML trees. While our current benchmarks focus on flat tables, the primary challenges we address stem from this embedded unstructured content. These challenges often require semantic inference, an area where traditional SQL struggles and large language models (LLMs) are critical in bridging the gap. Examples include inconsistencies in formatting, reliance on domain-specific knowledge, and implicit contextual cues necessary for accurate query interpretation.
Additionally, our approach is easily extendable to hierarchical data formats (e.g., HTML, nested JSON, HiTab) through structure decoding. These formats can be transformed into a flattened, normalized table representation and then processed using Weaver . However, as we discuss later, such datasets typically lack the complex hybrid queries involving multi-step reasoning and semantic inference, which are the primary focus of our work.
Filtering Hybrid Queries using Binder.
To filter hybrid queries from their original dataset counterparts, we leverage Binder’s query structure. Specifically, we identify queries that invoke any UDF, for example ‘QA‘ function as seen in the example below, which prompts the LLM with a yes/no question related to a single column. This serves as a reliable indicator that the query requires semantic reasoning beyond SQL capabilities. The example below demonstrates such a pattern, where the QA function is applied to assess contextual understanding. This filtering process ensures that the LLM is used for its intended purpose, semantic inference or interpretation, aligned with the methodology outlined in the "Our Approach" section.
Listing 24: Example NeuralSQL query using QA function
@@ Appendix E Comparison with Existing Datasets
Several benchmark datasets—such as Spider, HiTab, and BIRD—have contributed significantly to multi-hop and multi-table question answering. However, these benchmarks primarily emphasize symbolic SQL reasoning and do not align with the hybrid reasoning focus of Weaver , which requires coordinated symbolic (SQL-based) and semantic (LLM-based) reasoning.
@@ E.1 Evaluation on BIRD Dataset
We evaluate Weaver  on the BIRD benchmark using GPT-4o as the backend. Weaver  achieves an accuracy of 91%, solving nearly all tasks using symbolic SQL execution alone, without invoking any LLM-based semantic reasoning. These results demonstrate that while Weaver  is highly effective on BIRD, the dataset’s symbolic nature limits its value for testing hybrid reasoning capabilities. Most BIRD tasks can be completed in 3–4 SQL steps involving standard operations like joins, filtering, and sorting. For instance, consider the query:
Listing 25: Example execution plan
Question: Please list the lowest ten eligible free rates for students aged 5-17 in continuation schools. Eligible free rates Free Meal Count (Ages 5-17) / Enrollment (Ages 5-17)
@@
SQL: Join the frpm table with the schools table using School Type and SOCType, filtering for "Continuation High Schools".
SQL: Compute eligible free rates per school using the formula: Free Meal Count (Ages 5-17) / Enrollment (Ages 5-17).
@@
Step 1: SELECT f.* FROM frpm f JOIN schools s ON f.School Type = s.SOCType WHERE s.SOCType = Continuation High Schools’;
Step 2: CREATE TABLE eligible_free_rates AS SELECT School Name, School Type, Free Meal Count (Ages 5-17), Enrollment (Ages 5-17), (Free Meal Count (Ages 5-17) / Enrollment (Ages 5-17)) AS eligible_free_rate FROM continuation_schools;
Step3: CREATE TABLE sorted_eligible_free_rates AS SELECT * FROM eligible_free_rates ORDER BY eligible_free_rate ASC;
Step 4: CREATE TABLE lowest_ten_eligible_free_rates AS SELECT * FROM sorted_eligible_free_rates LIMIT 10;
@@
While Spider, HiTab, and BIRD advance multi-hop QA in important ways, none capture the core challenges of hybrid queries where symbolic (SQL) and semantic (LLM) steps must be interleaved. Weaver  is explicitly designed for such hybrid workflows, requiring intelligent planning, decomposition, and alternating execution paths—capabilities not required in these prior datasets.
@@ E.2 Execution Strategy in the Weaver Pipeline
Unlike methods like H-STAR, which execute SQL and LLMs in parallel, Weaver  selectively chooses whether SQL or LLM should handle each query part. This planning-based approach improves efficiency by avoiding the computational cost of parallel execution. As shown in Table 5: Number of API Calls Comparison per TableQA, H-STAR requires 8 LLM API calls per query, whereas Weaver  averages only 5.4, significantly reducing overhead and improving overall performance.

### v2

@@ Appendix B Dataset Details
- Short-Form Answering (WikiTQ): WikiTQ is a dataset designed for short-form answering where the steps to reach the answer can be relatively complex and the expected answer is a short piece of information. For example, the query “Which country had the most competitors?” in Figure 1. This task is ideal for testing how well information retrieval from a semi-structured table can be handled.
- Fact-Checking (TabFact): TabFact focuses on fact-checking tasks where the query involves verifying whether a particular statement is true or false. For example, the query, “Is the GDP of Japan in 2022 greater than that of Germany?” evaluates the framework’s ability to correctly interpret data points in complex tables and make valid judgments.
- Numerical Reasoning (FinQA): FinQA (Financial QA) is a dataset that requires the model to perform arithmetic operations or infer relationships between numerical values across different columns. For example, “What is the total revenue of Company X in 2021 after deducting expenses?” tests the model’s ability to handle numerical data and apply operations such as summation, subtraction, or other mathematical reasoning tasks.
@@
[…]
@@ Appendix C Few-Shot Prompting for Plan Generation
We use a few-shot prompting approach to generate plans during the planning stage. Through our analysis of various complex data problems, we identified three broad categories of transformation challenges that commonly arise. For each category, we manually crafted a representative example, rather than using examples from any particular dataset, to serve as in-context prompts. This was done to avoid memorization bias and to encourage cross-dataset generalization. These examples were specifically designed to capture the reasoning skills required for hybrid queries. We will include them in the appendix for clarity and reproducibility in the final version. Below are the three categories and the corresponding examples with sample tables:
@@
Listing 20: Few-shot examples for semantic reasoning from textual content
@@
"Indulge your senses with this botanical blend of rosemary and lavender. Gently
cleanses while nourishing your hair, leaving it soft, shiny, and revitalized."
7.99 4.99
@@
Listing 21: Few-shot examples for commonsense or background knowledge inference
@@
Question: Which location had the maximum time taken between dispatch at one location
and arrival or delivery at a subsequent location?
@@
Step 3: SQL - Pair Dispatched events with the corresponding Arrived at Hub or
Delivered events for each order/product and calculate the time difference.
Step 4: SQL - Display the final output with paired events, durations, and relevant
information sorted by Order_ID and Product.
@@
Listing 22: Few-shot examples for pre-processing or normalization of non-SQL-friendly columns
@@
Step 1: LLM - Format the column Placing by extracting only numerical values (e.g. 5
from Semifinal (5th)) and converting the text into numbers (e.g. Semifinal to 5,
12th to 12, 19 to 19).
Step 2: SQL - Retrieve the highest placing (rank) from the placing column by
selecting the minimum number in the list as lower number corresponds to higher rank.
Appendix D Additional Discussion
@@
Our core contribution lies in how we effectively leverage LLMs beyond what traditional SQL systems can offer, specifically in tasks involving column transformation and semantic inference. Weaver integrates LLM reasoning for two primary purposes: (i) handling tasks that SQL cannot perform, such as entity extraction or parsing ambiguous formats A.2 and (ii) inferring implicit knowledge based on pretraining, where the LLM is tasked with verifying a plan that involves semantic understanding Figure 2.
A detailed example from the WikiTQ dataset in Appendix A.2 (Listing 12) demonstrates the LLM’s capability to normalize complex date strings like Spring 1932 or 1935/36 into standardized YYYY formats. This transformation, impossible through SQL alone, was completed through a single LLM step with high consistency.
Listing 23: Example explaination
@@
Listing 24: Example explaination
Question: "What movies suitable for kids with positive reviews should be recommended
based on their reviews after 2018?"
@@
LLM: Utilize a Large Language Model (LLM) to evaluate whether a movie is suitable
for children by processing:
@@
Beyond Semi-structured Text:
We define semi-structured tables following prior work Gupta et al. (2020), focusing on unstructured or free-form content embedded within structured table formats, such as textual cells requiring semantic interpretation, rather than on hierarchical structures like JSON or HTML trees. While our current benchmarks focus on flat tables, the primary challenges we address stem from this embedded unstructured content. These challenges often require semantic inference, an area where traditional SQL struggles and large language models (LLMs) are critical in bridging the gap. Examples include inconsistencies in formatting, reliance on domain-specific knowledge, and implicit contextual cues necessary for accurate query interpretation.
Additionally, our approach is easily extendable to hierarchical data formats (e.g., HTML, nested JSON) through structure decoding. These formats can be transformed into a flattened, normalized table representation and then processed using Weaver . However, as we discuss later, such datasets typically lack the complex hybrid queries involving multi-step reasoning and semantic inference, which are the primary focus of our work.
Filtering Hybrid Queries using Binder:
To filter hybrid queries from their original dataset counterparts, we leverage Binder’s query structure. Specifically, we identify queries that invoke any UDF, for example ‘QA‘ function as seen in the example below, which prompts the LLM with a yes/no question related to a single column. This serves as a reliable indicator that the query requires semantic reasoning beyond SQL capabilities. The example below demonstrates such a pattern, where the QA function is applied to assess contextual understanding. This filtering process ensures that the LLM is used for its intended purpose, semantic inference or interpretation, aligned with the methodology outlined in the section 2.
Listing 25: Example Binder-style UDF SQL query using QA function
@@ Appendix E Comparison with Existing Datasets
Several benchmark datasets, such as Spider Yu et al. (2018), HiTab Cheng et al. (2022), and BIRD Li et al. (2023) have contributed significantly to multi-hop and multi-table question answering. However, these benchmarks primarily emphasize symbolic SQL reasoning and do not align with the hybrid reasoning focus of Weaver , which requires coordinated symbolic (SQL-based) and semantic (LLM-based) reasoning.
@@ E.1 Evaluation on BIRD Dataset
We evaluate Weaver on the BIRD benchmark using GPT-4o as the backend. Weaver achieves an accuracy of 91%, on a manually curated subset of 30 queries from the BIRD development set. We found that these sampled queries could be solved entirely using symbolic SQL execution alone, without invoking any LLM-based semantic reasoning. These results demonstrate that while Weaver is highly effective on BIRD, the dataset’s symbolic nature limits its value for testing hybrid reasoning capabilities. Most BIRD tasks can be completed in 3–4 SQL steps involving standard operations like joins, filtering, and sorting. For instance, consider the query:
Listing 26: Example execution plan
Question: Please list the lowest ten eligible free rates for students aged 5-17
in continuation schools. Eligible free rates Free Meal Count (Ages 5-17) /
Enrollment (Ages 5-17)
@@
SQL: Join the frpm table with the schools table using School Type and SOCType,
filtering for "Continuation High Schools".
SQL: Compute eligible free rates per school using the formula: Free Meal Count
(Ages 5-17) / Enrollment (Ages 5-17).
@@
Step 1: SELECT f.* FROM frpm f JOIN schools s ON f.School Type = s.SOCType WHERE
s.SOCType = Continuation High Schools’;
Step 2: CREATE TABLE eligible_free_rates AS SELECT School Name, School Type, Free
Meal Count (Ages 5-17), Enrollment (Ages 5-17), (Free Meal Count (Ages 5-17) /
Enrollment (Ages 5-17)) AS eligible_free_rate FROM continuation_schools;
Step3: CREATE TABLE sorted_eligible_free_rates AS SELECT * FROM eligible_free_rates
ORDER BY eligible_free_rate ASC;
Step 4: CREATE TABLE lowest_ten_eligible_free_rates AS SELECT * FROM
sorted_eligible_free_rates LIMIT 10;
@@
While Spider, HiTab, and BIRD advance multi-hop QA in important ways, none capture the core challenges of hybrid queries where symbolic (SQL) and semantic (LLM) steps must be interleaved. Weaver is explicitly designed for such hybrid workflows, requiring intelligent planning, decomposition, and alternating execution paths capabilities not required in these prior datasets.
@@ E.2 Execution Strategy in the Weaver Pipeline
Unlike methods like H-STAR, which execute SQL and LLM in parallel, Weaver  selectively chooses whether SQL or LLM should handle each query part. This planning-based approach improves efficiency by avoiding the computational cost of parallel execution. As shown in Table 5: Number of API Calls Comparison per Table QA, H-STAR requires 8 LLM API calls per query, whereas Weaver  averages only 5.4, significantly reducing overhead and improving overall performance.
@@ E.3 Rationale for Using SQL over Python
Appendix F Comparison with Related Table Reasoning Methods
Chain-of-Table Wang et al. (2024) represents an approach to table reasoning using iterative, chain-of-thought style planning. Weaver differs fundamentally in its methodology: it first generates a complete, high-level plan based on the question, table, and metadata, and only then begins executing each step. This global planning phase ensures interpretability and consistency across steps, similar to thinking several moves ahead in chess, rather than incrementally generating one step at a time. Because Chain-of-Table and ReAcTable both follow an incremental, step-by-step execution pattern, we included ReAcTable as a baseline. While ReAcTable captures a similar iterative reasoning paradigm, Weaver ’s novelty lies in decoupling reasoning from execution through its upfront, verifiable planning.
The TAG framework requires manually crafted code or logic for each user query and does not support direct natural language query execution without human intervention. Comparing such methods to natural-language-driven systems like Weaver would introduce unfair bias. TAG remains a relevant and valuable framework and will serve as a future benchmark once Weaver is extended to handle multi-table logic.
Appendix G Future Work
The modular design of Weaver naturally supports extensions to more complex settings. While our experiments focus on single-table English datasets to isolate hybrid reasoning capabilities, the model is not inherently limited to such settings. Preliminary tests on more complex datasets support this generalization: on BIRD, a multi-table benchmark, primarily solving tasks via only SQL, and on HiTab, which contains hierarchical tables, flattening and normalizing the schema enabled similar SQL-based execution. These results demonstrate that Weaver can adapt to multi-table and hierarchical structures when semantic inference is not required (see Appendix E).
Looking forward, Weaver can be extended along three directions. First, hierarchical or nested tables could be supported by preserving context across nested attributes and extending the LLM+SQL planning strategy. Second, multi-table datasets can be handled through schema-aware planning and cross-table reasoning while maintaining hybrid execution. Third, non-English tables can be incorporated by leveraging multilingual LLMs or language-specific prompts. These directions highlight Weaver ’s potential to scale seamlessly to diverse, real-world, and multilingual datasets, extending its applicability well beyond the single English tables evaluated here.